How many grams of tetraphosphorus decaoxide are required for the complete reaction of 21.6 grams of perchloric acid?
1
Expert's answer
2017-12-08T07:00:53-0500
P4O10 + 4HClO4 = 4HPO3 + 2Cl2O7
m(HClO4) = 21.6 g
Mr(HClO4) = 1 + 35.5 + 16*4 = 100.5 g/mol
Mr(P4O10) = 31*4 + 16*10 = 284 g/mol
m(P4O10) = 284*21.6/100.5 = 61.0 g